OpenStudy (anonymous):

What is the slope of the line between (3, -4) and (-2, 1)? 1 2 -2 -1

OpenStudy (anonymous):

isn't there an equation to find this besides plotting the point on a chart?

OpenStudy (skullpatrol):

Yes. What does "slope" mean?

OpenStudy (anonymous):

slope would be well the angle of the slope in which a line travels across a cordinate plane?

OpenStudy (anonymous):

slope=(y2-y1)/(x2-x1) =(1-(-4))/(-2-3) =(1+4)/(-5) =5/-5 =-1

OpenStudy (skullpatrol):

Why are you giving the answer away? @madrockz ???

OpenStudy (anonymous):

I think I remember doing that formula a while back

OpenStudy (anonymous):

wat is wrong in dat?

OpenStudy (anonymous):

(3, -4) and (-2, 1) so... (1 -(- 4) / (-2 - 3) 1 - (-4) = 5... -2 - 3 = -5... 5/-5 = -1 :D

OpenStudy (skullpatrol):

@madrockz is the student learning anything?

OpenStudy (anonymous):

if the student wish to learn...he will first understand the steps and find out where he was wrong and get it corrected.

OpenStudy (skullpatrol):

Failure to follow the guidelines may lead to your suspension or banning from the site. If you see any violations, please report them by using a report abuse link. OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

ohhh....ok u must have said dat in a polite manner

OpenStudy (anonymous):

I understand the formula now :)

OpenStudy (skullpatrol):

trial & error is not an efficient way of learning math

OpenStudy (anonymous):

Well it wasn't trial and error, i just forgot what the formula was, i completely understand the formula of (y2 - y1) / (x2 - x1) :)

OpenStudy (anonymous):

hey brad...have u understood my ans?i need to know ur opinion

OpenStudy (anonymous):

Yes I did

OpenStudy (anonymous):

ok thanks

OpenStudy (skullpatrol):

See if you understood what "slope" meant you would not have forgotten, right?

OpenStudy (skullpatrol):

Teaching MEANING is the most important thing...not formulas!

OpenStudy (anonymous):

I think I understand the meaning of a slope

OpenStudy (anonymous):

SWAG is watching me :o

OpenStudy (skullpatrol):

@Brad1996 what does it mean?

OpenStudy (anonymous):

To me slope is the angle of which a line travels on a cordinate plane. The slope is showed normally by a fraction or whole number

OpenStudy (skullpatrol):

OK, what is negative slope?

OpenStudy (anonymous):

a negative slope is a slope thats on a downwards path. Like the first cordinate is at 0,0 then the second one is at 5,-5

OpenStudy (skullpatrol):

What is the value for a vertical slope?

OpenStudy (anonymous):

x = 5 would be a vertical slope

OpenStudy (anonymous):

the x cordinate never changes and the line just goes straight up while the y value is free to change :)

OpenStudy (skullpatrol):

I asked you for the value of the slope.

OpenStudy (anonymous):

undefined is what your asking for?

OpenStudy (anonymous):

because I know that vertical lines slope is undefined and horizontal is 0

OpenStudy (skullpatrol):

Yes, undefined... also, what is the value of a horizontal slope?

OpenStudy (anonymous):

0

OpenStudy (skullpatrol):

But you have ONLY memorized them. Can you explain them to me?

OpenStudy (anonymous):

its 0 because a horizontal line is just going straight to the right or left. It is not sloping at any angle which would make its value of slope 0

OpenStudy (anonymous):

Vertical is undefined because its just going straight up or straight down and has well a undefined slope lol

OpenStudy (anonymous):

I really need to get working again, been on this question for half an hour lol

OpenStudy (anonymous):

thanks for the help though :)

OpenStudy (skullpatrol):

Thanks for trying to understand :-)

OpenStudy (anonymous):

lol I got plenty more questions your more then welcome to help me understand them :)